OpenStudy (anonymous):

how do I solve a three equation problem using elimanation. -5x+8y+3z=122 7x-2y-12z=-783 -8x-9y-10z=15 ? I just want to know how to solve step by step -_-

OpenStudy (anonymous):

Essentially the same way you solve a 2x2 linear system. It is a fairly time consuming and detail oriented proccess.

OpenStudy (anonymous):

A -5x+8y+3z=122 B 7x-2y-12z=-783 C -8x-9y-10z=15 pick 2 to solve for, then it spits out another (D) A B --- D Solve the other two, then it spits out another (E) B C --- E then solve for D and E. After that, x, y, z should all check in the original 3 equations.

OpenStudy (anonymous):

I hope that helps!

OpenStudy (anonymous):

and what am i suppose to solve after i get d e ? i get all big numbers

OpenStudy (anonymous):

for the top two i got 23x-45z=-3,010 and fot the bottom two i got 79x-88z=-7077 now im suppose to get rid of one letter but heres where im stuck at

OpenStudy (anonymous):

Your numbers are going to be different depending out what you used to eliminate. But it is good that you eliminated just one variable (y) between the two. I forgot to mention that. Now you take the two you got and eliminate another variable. (x or z). If you want to eliminate x, multiply -79 on the top and 23 on the bottom. Then then the x's will cancel out.

OpenStudy (anonymous):

okay so for the top i got 1817x-3555z=-237790 and the bottom is 1817x-2024=-162771 then i work it out and i get -5579z=-400561 if i divide z to both sides i get a decimal am i suppose to get that or a whole number ?

OpenStudy (anonymous):

Our numbers will be different, guaranteed. There are an infinite amount of ways to solve it. I tried running through it... but I'm not great with details. So, my answer didn't add up. But this is basically how you are supposed to do it: A -5x+8y+3z=122 B 7x-2y-12z=-783 C -8x-9y-10z=15 pick 2 to solve for, then it spits out another (D) 7* A -5x+8y+3z=122 -35x+56y+21z=854 5* B 7x-2y-12z=-783 35x-10y-60z=-3915 --- D 46y-39z=-3061 Solve the other two, then it spits out another (E) 8* B 7x-2y-12z=-783 56x-16y-96z=-6264 7* C -8x-9y-10z=15 -56x-63y-70z=105 --- E -79y-166z=-6159 then solve for D and E. 79* D 46y-39z=-3061 3634y-3081z=-241819 46* E -79y-166z=-6159 -3634y-7636z=-283314 --- -10717z=-525133 /-10717 /-10717 z=49 Repeat, but solve for y: -166* D 46y-39z=-3061 -7636y+6474z=508126 39* E -79y-166z=-6159 -3081y-6474z=-240201 --- -10717y=267925 /-10717 /-10717 y=-25 (z=49, y=-25) C -8x-9(-25)-10(49)=15 -8x+225-490=150 -8x-265=150 +265 +265 -8x=-115 /-8 /-8 x=14.375 (z=49, y=-25, x=14.375) After that, x, y, z should all check in the original 3 equations... which mine didn't... Sorry :(

OpenStudy (anonymous):

OK OK OK, I got. I don't know where I messed up early, but this is the solution. Very detailed oriented, as mentioned. Hope this helps you out! A -5x+8y+3z=122 B 7x-2y-12z=-783 C -8x-9y-10z=15 pick 2 to solve for, then it spits out another (D) 7* A -5x+8y+3z=122 -35x+56y+21z=854 5* B 7x-2y-12z=-783 35x-10y-60z=-3915 --- D 46y-39z=-3061 Solve the other two, then it spits out another (E) 8* B 7x-2y-12z=-783 56x-16y-96z=-6264 7* C -8x-9y-10z=15 -56x-63y-70z=105 --- E -79y-166z=-6159 then solve for D and E. 79* D 46y-39z=-3061 3634y-3081z=-241819 46* E -79y-166z=-6159 -3634y-7636z=-283314 --- -10717z=-525133 /-10717 /-10717 z=49 Repeat, but solve for y: -166* D 46y-39z=-3061 -7636y+6474z=508126 39* E -79y-166z=-6159 -3081y-6474z=-240201 --- -10717y=267925 /-10717 /-10717 y=-25 (z=49, y=-25) A -5x+8(-25)+3(49)=122 -5x-200+147=122 -5x-53=122 +53 +53 -5x=175 /-5 /-5 x=-35 (z=49, y=-25, x=-35) After that, x, y, z should all check in the original 3 equations. A -5(-35)+8(-25)+3(49)=122 (Checks) B 7(-35)-2(-25)-12(49)=-783 (Checks) C -8(-35)-9(-25)-10(49)=15 (Checks!)

OpenStudy (anonymous):

thank you for helping me guys i appericate it and i understand now :)